OpenStudy (anonymous):

Cot^2x/cscx+1=cscx-1

OpenStudy (anonymous):

\[\frac{ \cot^2(x) }{ \csc(x)+1 } = \csc(x) -1 \] If this correct?

OpenStudy (anonymous):

Do you want to prove it, or find the x's which make it true?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Prove it

OpenStudy (anonymous):

cot(x) = cos(x)/sin(x) therefore: cot^2(x) = (cosx/sinx)(cosx/sinx) csc(x) = (1/sinx)

OpenStudy (anonymous):

Thank you.